OpenStudy (lxelle):

The equation of a curve is sqrt x + sqrt y = sqrt a, where a is a positive constant. Express dy/dx in terms of x and y.

OpenStudy (adi3):

@mim_3 any idea

OpenStudy (adi3):

@Mimi_x3

OpenStudy (nikvist):

\[\sqrt{x}+\sqrt{y}=\sqrt{a}\]\[\frac{1}{2\sqrt{x}}+\frac{1}{2\sqrt{y}}\frac{dy}{dx}=0\]\[\frac{dy}{dx}=-\sqrt{\frac{y}{x}}\]

OpenStudy (lxelle):

Pls explain your working then?

OpenStudy (adi3):

I m back

OpenStudy (adi3):

sorry for late reply lost connection

OpenStudy (lxelle):

@Adi3 you weren't even helping me -.-

OpenStudy (jhannybean):

Looks like this is a separable differential equation. So in terms of finding dy/dx in terms of y and x, we can take the derivative of the entire function with respect to x, and then just solve for dy/dx. That's my take on it.

OpenStudy (adi3):

i always loose connection

OpenStudy (adi3):

whenever I trie to reply no conction

OpenStudy (lxelle):

Implicit equation tho.

OpenStudy (jhannybean):

Oh yes, that is correct. Implicit as we are solving for \(y'\).

OpenStudy (adi3):

I don't know how to solve this. ohh shoot don't know why

OpenStudy (lxelle):

@Adi3 Get lost pls. I don't need your help.

OpenStudy (jhannybean):

\[\sf \sqrt{x} +\sqrt{y} = c\]\[d(\sqrt{x}+\sqrt{y}=c)\]\[d(\sqrt{x})+d(\sqrt{y}) = d(c)\]\[\frac{1}{2x^{1/2}} +\frac{1}{2y^{1/2}} \cdot y' = 0\]Solving for \(y'\) we get \[\frac{1}{2y^{1/2}}y' =-\frac{1}{2x^{1/2}}\]\[\boxed{y'=-\dfrac{y^{1/2}}{x^{1/2}} = -\left(\dfrac{y}{x}\right)^{1/2} = -\sqrt{\dfrac{y}{x}}} \]

OpenStudy (jhannybean):

I labeled \(\sqrt{a}\) as c, indicating that it is a constant.

OpenStudy (lxelle):

the d is referring to dy/dx isit

OpenStudy (jhannybean):

\[d=\frac{d}{dx}\] And thus implicitly differentiating y with respect to x, we have \(\dfrac{dy}{dx}\)

OpenStudy (jhannybean):

Does that make sense?

OpenStudy (lxelle):

YESSS. Thanks so much.

OpenStudy (jhannybean):

Awesome.